# Uncovering the Hidden Threat of Text Watermarking from Users with Cross-Lingual Knowledge

### v1

@@ 6 Results
We perform experiments to answer the following research questions:
RQ3: How do different adversarial approaches to cross-lingual watermark evasion perform? Specifically, how do attacks fare under our novel asymmetrical threat model, where the output language differs from the pivotal language?
@@ 1.6 RQ1: Watermarking Performance under Cross-lingual Setting
Figure 2 shows the result of obtaining the decreases of each criterion score after watermarking, indicating the text coherency is the most affected aspect in all languages. Moreover, XSIR and EXP showcase the largest degrade of the quality. Our analysis indicates that GPT-Judger frequently flags issues related to clarity of writing, including repetition and the overlapping of different language characters or unknown symbols.
@@
(a) KGW
(b) Unigram
(c) XSIR
(d) EXP
To further analyze the GPT-Judger results, we calculate soft-win rates in Table 1. A soft-win occurs when the judge favors the watermarked text or issues a tie. Among all methods, KGW achieves the highest soft-win rate. Nonetheless, the GPT-Judger occasionally struggles to deliver definitive judgments, particularly for non-English texts, as indicated by the high tie rates compared to English. To objectively capture this impact, we introduce our new metric rooted in the Self-BLEU (SB) metric.
2 : SelfBleu results for all watermark methods. For KGW, Unigram and XSIR, $\gamma=0.5$ and $\delta=2.0$. We consider this setting is the best for these methods with different languages. More results with varying the $\gamma$ and $\delta$ values in are in Appendix
@@
4 : Watermark detection ROC curves with AUC after attacks. We fix $\gamma=0.5$ and $\delta=2.0$ for KGW, Unigram, and XSIR. For all attacks, the watermark threshold is calculated automatically by comparing unwatermarked and attacked watermarked text score for $100$ generations. Left: we perform translation attacks on KGW, Unigram, and EXP. Right: Only the leftmost column represent asymmetrical translation attack setting. We apply more invasive attacks to XSIR to highlight its robustness
@@
Diversity. Previous research has used the Self-BLEU (SB) metric to assess the diversity of watermarked text (Dathathri et al., 2024). Table 2 includes results from SB and the Adjusted Diversity (AD) metrics for various watermark methods. While SB suggests high diversity, especially for Chinese and Indonesian, it can be misleading; low SB scores (e.g., $0.04$ for Chinese) might indicate text degradation, as shown by higher AD scores ($\geq 0.44$). This pattern holds across other languages. When we employ larger $\delta$ values with $\gamma=0.5$, AD ratios are higher compared to SB, indicating unrealistic diversity in watermarked text. More detailed results with different hyper-parameters are in Table 3 in Appendix A(5)(5)(5)Language mixing can be due to combined-language prompts, model size, or vocabulary coverage, and watermarking exacerbate the problem.
@@
Detectability. Our detection analysis examines both performance without attack and attack resilience across languages. Figures 3 and 4 show the ROC curves for all watermarking methods before and after attacks. Prior to any attacks, all methods demonstrate strong detectability. For KGW, Unigram, and XSIR, we identified $\delta=2.0$ and $\gamma=0.5$ as the optimal parameters yielding the best detection results. However, the role of employing different hyper-parameters for detection differs from that for text quality. When larger values of $\gamma$ are used with $delta=2.0$, detection scores are adversely affected, especially for Unigram method. This is counterintuitive to the green/red splitting methodology as in KGW and Unigram since using $delta=2.0$ with smaller $\gamma=0.1$ yield higher detection scores. We show detailed results in Figure 7 in which we illustrate the close relationship between $\gamma$ and $\delta$ for methods like KGW and Unigram in Appendix B.
2.6 RQ2: Watermarking Resiliencies to Translation Attacks.
@@
Regarding attack resilience, all methods vulnerable to our proposed asymmetrical attacks. In Figure 4 (left), we perform asymmetrical translation attacks against KGW, Unigram, and EXP. The results showcase the watermark detectability under translation attacks varies by language Ghanim et al. (2024). For instance, KGW shows the lowest AUC of 0.55 for Arabic-English, while Unigram presents a worst-case AUC of 0.57 for English-Arabic and Arabic-English. In the EXP method as shown in Figure 4 (c), all attacks are notably invasive, likely due to its larger hashing window $4$ compared to KGW’s recommended window of $1$.
@@
In Figure 4 (right), we evaluate all of our asymmetrical attacks to the more robust XSIR method. Although XSIR was developed for symmetrical attacks, it remains susceptible to asymmetrical attacks. This vulnerability underscores the severity of attack scenarios where the pivot language differs from the output language. XSIR’s semantic clusters vary across languages, weakening its watermark robustness when the pivot and output languages differ. In contrast, back-translate attacks preserve the watermark better, as translating back to the pivot language realigns the semantic clusters as shown in Figure 4 (f).
@@
Additionally, KGW and EXP have more predictable results than XSIR and Unigram where different languages affect the watermark signal differently. For instance, detection outcomes for translation and translation-paraphrase attacks remain consistently tight across languages for KGW and EXP, whereas XSIR and Unigram exhibit greater variability, with some languages deviating significantly. This behavior may stem from technical challenges such as preserving text length during translation.
@@ 3.6 RQ3: Asymmetrical Translation Attacks Raise a New Threat
Despite XSIR being specifically designed for cross-lingual scenarios, our results shown in Figure 4 demonstrate that it remains vulnerable to post-generation translation attacks. Particularly, the vulnerability is pronounced for Chinese and Arabic languages, suggesting that XSIR’s effectiveness varies significantly by language. This contrasts with XSIR proposed adversarial scenario of translating prompts both before and after generation, indicating that post-generation attacks can be more effective at evading detection.
Our analysis reveals a critical asymmetry: translations TO English yield better detection rates than translations FROM English. This asymmetry appears to be rooted in XSIR’s semantic clustering characteristics, as we found that XSIR semantic clusters have more similar English words than other languages. This disparity is likely due to XSIR’s semantic clustering, which groups English words more consistently than words from other languages. Consequently, XSIR appears to perform more robustly when English is the target language rather than the source language.
These findings highlight the need for more robust cross-lingual watermarking approaches that can maintain effectiveness across different languages and resist sophisticated attack pipelines.
@@ 7 Conclusion
In this paper, we assess four distinct watermarking techniques across four different languages, considering realistic watermark removal attacks in a cross-lingual context. We examine the effectiveness of these watermarking methods by evaluating the quality of the watermarked text and the resilience of the watermark signal following such attacks. Our findings indicate that various watermarking techniques respond differently to language translations when the watermarked text is modified. We propose that there is a need for more robust cross-lingual watermarking solutions that integrate the predictability of KGW and EXP across languages with the more resilient approaches of Unigram and XSIR.
8 Limitations and Discussions
In this study, we evaluate four distinct watermarking methods across four languages, focusing on practical removal attacks in a cross-lingual context. We believe that more research should include languages with rich syntax and structure and suggest exploring datasets beyond C4, given its generalized nature.
The attacks presented are invasive, significantly weakening the watermark signal, but can be countered through various strategies. We believe that developing more resilient cross-lingual watermarking techniques should integrate KGW and EXP predictability with the robust Unigram method. XSIR could be instrumental here, as it manipulates text semantics over syntax. We first analyze why XSIR falls short against translation attacks and then explore potential solutions.
Our investigation into XSIR revealed that its shortcomings in handling translation attacks are largely technical. XSIR clusters semantically similar words across languages by constructing a graph from a comprehensive dictionary using text transformations. It identifies connected components (CCs), but this process can be problematic for languages with diverse character sets. For instance, CCs rely on exact matches, which may prevent clustering of semantically similar words.
To enhance XSIR’s robustness, we suggest modifying the detection process. Given that English handles XSIR cluster transformations effectively, translating non-English outputs to English before detection could ensure watermark signal identification. This semantic-focused approach can extend to other methods, such as PostMark  (Chang et al., 2024).
Additionally, XSIR’s robustness can be improved by refining the use of CCs. Changing node content to latent word representations, formulated as numerical vectors, would enable more meaningful exact matching, enhancing clustering accuracy across different languages.
@@ 9 Ethics Consideration
  • Cai et al. (2024) Kunbei Cai, Zhenkai Zhang, Qian Lou, and Fan Yao. 2024. Wbp: Training-time backdoor attacks through hardware-based weight bit poisoning. In European Conference on Computer Vision, pages 179–197. Springer.
@@
  • Lou et al. (2022) Qian Lou, Yepeng Liu, and Bo Feng. 2022. Trojtext: Test-time invisible textual trojan insertion. In The Eleventh International Conference on Learning Representations.
@@
  • Singh and Zou (2023) Karanpartap Singh and James Zou. 2023. New evaluation metrics capture quality degradation due to llm watermarking. Transactions on Machine Learning Research.
@@
  • Xue et al. (2024a) Jiaqi Xue, Qian Lou, and Mengxin Zheng. 2024a. Badfair: Backdoored fairness attacks with group-conditioned triggers. In Findings of the Association for Computational Linguistics: EMNLP 2024, pages 8257–8270.
  • Xue et al. (2024b) Jiaqi Xue, Mengxin Zheng, Yebowen Hu, Fei Liu, Xun Chen, and Qian Lou. 2024b. Badrag: Identifying vulnerabilities in retrieval augmented generation of large language models. arXiv preprint arXiv:2406.00083.
  • Xue et al. (2024c) Jiaqi Xue, Mengxin Zheng, Ting Hua, Yilin Shen, Yepeng Liu, Ladislau Bölöni, and Qian Lou. 2024c. Trojllm: A black-box trojan prompt attack on large language models. Advances in Neural Information Processing Systems, 36.
  • Zhang et al. (2025) Hang Zhang, Qian Lou, and Yanshan Wang. 2025. Towards safe ai clinicians: A comprehensive study on large language model jailbreaking in healthcare. arXiv preprint arXiv:2501.18632.
@@
  • Zheng et al. (2023) Lianmin Zheng, Wei-Lin Chiang, Ying Sheng, Siyuan Zhuang, Zhanghao Wu, Yonghao Zhuang, Zi Lin, Zhuohan Li, Dacheng Li, Eric Xing, et al. 2023. Judging llm-as-a-judge with mt-bench and chatbot arena. Advances in Neural Information Processing Systems, 36:46595–46623.

### v2

@@ 1.6 RQ1: Watermarking Performance under Cross-lingual Setting
In Figures 2 and  3 we show the result of obtaining the decrease percentage in quality after watermarking. In these figures, we only present the coherency criterion since its scores reflect the most affected criterion in all languages according to GPT-Judger. The relative decrease in the figure is calculated as:
$(\frac{\text{unwatermarked\_score}-\text{watermarked\_score}}{\text{unwatermarked\_score}})100$ )6(
We use C4 to reflect a more generalized text completion scenario, and we use LFQA for low-entropy instruction-following completions. The relative impact of coherency decrease varies between languages in the same watermarking method, particularly for C4 dataset. While Chinese and Indonesian often show higher coherency decrease than English and Arabic in several methods, this impact is reduced in low-entropy completions. While this suggests easier watermarking for instruction-following scenarios, it can affect the detectability of the watermark as shown in Table 7 in the appendices. Among non-English languages, Arabic seems to be easier to watermark under KGW and Unigram methods, but more difficult to watermark under XSIR. In terms of watermarking method, KGW has minimal impact of quality for both datasets, while XSIR and EXP showcase the largest degrade of the quality across all languages. In Appendix  A, Figure 7, we show more results for additional languages.
@@
To further analyze the GPT-Judger results, we calculate soft-win rates in Table 1. Among all methods, KGW achieves the highest soft-win rate. Nonetheless, the GPT-Judger occasionally struggles to deliver definitive judgments, particularly for non-English texts, as indicated by the high tie rates.
3 : GPT-judge coherency criterion results for Cohere model on LFQA dataset, computed as the average of watermarked and unwatermarked scores for $500$ generations.
LLM Judge Language Bias. Our fairness experiments reveal significant language biases in GPT-Judger’s final verdict, which may explain some of the result in Figure 2. The results of these experiments are presented in Appendix  C Tables  9 and  10. When evaluating identical content translated into different languages (Translation Experiment), we found clear preferences for certain languages regardless of content quality. German was most preferred, followed by Arabic and English, while Chinese, Persian, and Indonesian received consistently lower scores. In a complementary ”Paraphrase Experiment” using same-language text pairs, we observed judges showing higher TIE rates for some languages over the others. Additionally, the judger consistently favored paraphrased texts generated by their own model family over original texts. These biases suggest that the high TIE rates in our watermarking evaluations likely affected by the biased preferences we saw in the ”translation experiment” rather than actual quality equivalence judgments. For detailed analysis including position bias investigation and cross-language comparison metrics, see Appendix C. These findings align with token bias studies by Zheng et al. and model self-preference observations by Ye et al. (2024), and therefore, llm-as-a-judge paradigms should be coupled with some sort of debiasing procedures to ensure fair choice between texts of different languages.
2 : SelfBleu results for all watermark methods. For KGW, Unigram and XSIR, $\gamma=0.5$ and $\delta=2.0$. We consider this setting is the best for these methods with different languages. More results with varying the $\gamma$ and $\delta$ values in are in Appendix.
@@
4 : Watermark detection ROC curves with AUC after attacks. For all attacks, the watermark threshold is calculated automatically by comparing unwatermarked and attacked watermarked text score for $100$ generations. Left: we perform symmetrical translation attacks on KGW, Unigram, and EXP. Right: Only the leftmost column represent symmetrical translation attack setting. We further apply more invasive attacks to XSIR to highlight its robustness.
@@
Diversity. To more objectively quantify quality of text with a predefined text quality metric, we introduce our new metric rooted in the Self-BLEU (SB) metric. In our experiments, the Adjusted Diversity (AD) is used to further assess the quality of watermarked text along with the assessment provided by the GPT-Judger’s coherency scores. Previous research has used the SB metric to assess the diversity of watermarked text (Dathathri et al., 2024). Table 2 includes results from SB and the Adjusted Diversity (AD) metrics for various watermark methods. While SB suggests high diversity for Chinese and Indonesian, it can be misleading; low SB scores (e.g., $0.04$ for Chinese) might indicate text degradation, as shown by higher AD scores ($\geq 0.44$). This pattern holds across other languages. When we employ larger $\delta$ values with $\gamma=0.5$, AD ratios are higher compared to SB, indicating unrealistic diversity in watermarked text. More detailed results with different hyper-parameters are in Table 3 in Appendix A.
@@
Detectability. Our detection analysis examines both performance without attack and attack resilience across languages. Prior to any attacks, all methods demonstrate strong detectability as shown in Figure 8 in Appendix B in which all methods across all languages achieve $\geq 0.99$ AUC scores. For KGW, Unigram, and XSIR, we identified $\delta=2.0$ and $\gamma=0.5$ as the optimal parameters yielding the best detection results. However, the role of employing different hyper-parameters for detection differs from that for text quality. When larger values of $\gamma$ are used with $delta=2.0$, detection scores are adversely affected, especially for Unigram method. We show detailed results in Figure 9 in which we illustrate the close relationship between $\gamma$ and $\delta$ for methods like KGW and Unigram in Appendix B.
2.6 RQ2: Watermarking Resilience to Translation Attacks.
@@
Regarding attack resilience, all methods are vulnerable to our attack pipeline. In Figure 4 (left), we perform asymmetrical translation attacks against KGW, Unigram, and EXP. The results showcase the watermark detectability under translation attacks varies by language. For instance, KGW shows the lowest AUC of $0.55$ for Arabic-English, while Unigram presents a worst-case AUC of $0.57$ for English-Arabic and Arabic-English. For EXP ( Figure 4 (c)), all attacks are notably invasive, likely due to its larger $k$-gram hashing window of $4$ compared to KGW’s recommended window of $1$.
@@
In Figure 4 (right), we evaluate all of our asymmetrical attacks to the more cross-lingual XSIR method. Although XSIR was developed for symmetrical attacks, it remains susceptible to asymmetrical attacks. This vulnerability underscores the severity of attack scenarios where the pivot language differs from the output language. XSIR’s semantic clusters vary across languages, weakening its watermark robustness when the pivot and output languages differ. In contrast, back-translate attacks preserve the watermark (Figure 4 (f)).
@@
Finally, KGW and EXP have more predictable results than XSIR and Unigram where different languages affect the watermark signal differently. For instance, detection outcomes for translation and translation-paraphrase attacks remain consistently tight across languages for KGW and EXP, whereas XSIR and Unigram exhibit greater variability, with some languages deviating significantly. This behavior may stem from technical challenges such as preserving text length during translation.
@@ 3.6 RQ3: Asymmetrical Translation Attacks Raise a New Threat
Despite XSIR being specifically designed for cross-lingual scenarios, our results shown in Figure 4 demonstrate that it remains vulnerable to post-generation translation attacks. Particularly, the vulnerability is pronounced for Chinese and Arabic languages, suggesting that XSIR’s effectiveness varies significantly by language. This contrasts with XSIR proposed adversarial scenario of translating prompts before generation and translating them back into the original language.
From Figures  4 and  11, our analysis reveals a critical asymmetry: starting the prompt with ENGLISH seems to yield better detection rates than starting with other languages. In other words, considering the interpolation of languages in the context of only using a language as pivotal to evade watermarking is not enough to assess the robustness of watermarks in cross-lingual manner.
The attacks presented are invasive, significantly weakening the watermark signal, but can be countered through various strategies. We believe that developing more resilient cross-lingual watermarking techniques should integrate KGW and EXP predictability with the robust Unigram method. XSIR could be instrumental here, as it manipulates text semantics over syntax.
Our investigation into XSIR revealed that its shortcomings in handling translation attacks are largely technical. XSIR clusters semantically similar words across languages by constructing a graph from a comprehensive dictionary that contains pairs of related words in multiple languages. It then creates connected components (CCs) to connect all similar nodes together through string matching. This creates a mapping that is used to broadcast the same bias to the logits of all related words in a cluster. However, the process in which the words in the dictionary are connected to a specific model’s vocabulary is tokenizer-dependent. This is because some tokenizers store vocabularies as unicode characters as a result of the encoding process of tokenization methods of different vocabularies. English tokens are usually stored as English alphabets, but this is not the case for non-English tokens. This creates clusters of matching English tokens only, bypassing similar words in other languages.
To enhance XSIR’s robustness, one can modify the detection process to translate a generated text to the original language of the prompt. Given that English handles XSIR cluster effectively, translating non-English outputs to English before detection could ensure watermark signal identification. However, this only works if the language in which the first prompt was generated is English. Clearly, this is not a viable option in a cross-lingual setting in which the prompt language can be non-English.
@@ 7 Conclusion
Our study reveals significant variations in LLM watermarking effectiveness across languages. KGW best preserves text quality, while all methods show vulnerability to translation attacks, especially asymmetrical ones where pivot and output languages differ. We identified substantial language biases in LLM-based evaluations, with GPT-Judger showing clear preferences for certain languages over others. Our findings demonstrate that current watermarking approaches inadequately address cross-lingual complexities, as effectiveness depends not only on the target language but also on specific language paths during attacks. This highlights the need for more robust cross-lingual watermarking methods that maintain effectiveness across diverse linguistic contexts.
8 Limitations
In this study, we evaluate four distinct watermarking methods across four languages—English, Arabic, Chinese, and Indonesian—focusing on practical quality evaluation and removal attacks in a cross-lingual context. We also extend the investigation to Turkish and Hindi using the CohereAI model. However, we believe that more research should include a wider variety of languages - both low-resource languages and those with complex syntactic and grammatical features. Additionally, we have included studies on C4 and LFQA datasets but we think expanding the datasets beyond these two could be beneficial in ensuring final insights about a language or a watermarking method.
@@ 9 Ethics Consideration
Acknowledgments
We would like to thank the reviewers for their helpful comments that strengthened this paper. We extend our thanks to King Khalid University for its support during the lifetime of this project.
@@ Acknowledgments
  • Al Ghanim et al. (2023) Mansour Al Ghanim, Muhammad Santriaji, Qian Lou, and Yan Solihin. 2023. Trojbits: A hardware aware inference-time attack on transformer-based language models. In ECAI 2023, pages 60–68. IOS Press.
@@
  • Cohere et al. (2025) Team Cohere, Arash Ahmadian, Marwan Ahmed, Jay Alammar, Milad Alizadeh, Yazeed Alnumay, Sophia Althammer, Arkady Arkhangorodsky, Viraat Aryabumi, Dennis Aumiller, et al. 2025. Command a: An enterprise-ready large language model. arXiv preprint arXiv:2504.00698.
@@
  • Koo et al. (2023) Ryan Koo, Minhwa Lee, Vipul Raheja, Jong Inn Park, Zae Myung Kim, and Dongyeop Kang. 2023. Benchmarking cognitive biases in large language models as evaluators. arXiv preprint arXiv:2309.17012.
  • Krishna et al. (2023) Kalpesh Krishna, Yixiao Song, Marzena Karpinska, John Wieting, and Mohit Iyyer. 2023. Paraphrasing evades detectors of ai-generated text, but retrieval is an effective defense. Advances in Neural Information Processing Systems, 36:27469–27500.
@@
  • Kurita et al. (2020) Keita Kurita, Paul Michel, and Graham Neubig. 2020. Weight poisoning attacks on pre-trained models. arXiv preprint arXiv:2004.06660.
@@
  • Lou et al. (2024) Qian Lou, Xin Liang, Jiaqi Xue, Yancheng Zhang, Rui Xie, and Mengxin Zheng. 2024. Cr-utp: Certified robustness against universal text perturbations on large language models. In Findings of the Association for Computational Linguistics: ACL 2024, pages 9863–9875.
@@
  • Pezeshkpour and Hruschka (2023) Pouya Pezeshkpour and Estevam Hruschka. 2023. Large language models sensitivity to the order of options in multiple-choice questions. arXiv preprint arXiv:2308.11483.
  • Qi et al. (2021) Fanchao Qi, Mukai Li, Yangyi Chen, Zhengyan Zhang, Zhiyuan Liu, Yasheng Wang, and Maosong Sun. 2021. Hidden killer: Invisible textual backdoor attacks with syntactic trigger. arXiv preprint arXiv:2105.12400.
@@
  • Xu et al. (2024) Jiashu Xu, Fei Wang, Mingyu Derek Ma, Pang Wei Koh, Chaowei Xiao, and Muhao Chen. 2024. Instructional fingerprinting of large language models. arXiv preprint arXiv:2401.12255.
  • Xue and Lou (2022) Jiaqi Xue and Qian Lou. 2022. Estas: Effective and stable trojan attacks in self-supervised encoders with one target unlabelled sample. arXiv preprint arXiv:2211.10908.
  • Xue et al. (2023a) Jiaqi Xue, Mengxin Zheng, Ting Hua, Yilin Shen, Yepeng Liu, Ladislau Bölöni, and Qian Lou. 2023a. Trojllm: A black-box trojan prompt attack on large language models. Advances in Neural Information Processing Systems, 36:65665–65677.
  • Xue et al. (2023b) Jiaqi Xue, Mengxin Zheng, Yi Sheng, Lei Yang, Qian Lou, and Lei Jiang. 2023b. Trojfair: Trojan fairness attacks. In Proceedings of the 1st ACM Workshop on Large AI Systems and Models with Privacy and Safety Analysis, pages 47–56.
  • Ye et al. (2024) Jiayi Ye, Yanbo Wang, Yue Huang, Dongping Chen, Qihui Zhang, Nuno Moniz, Tian Gao, Werner Geyer, Chao Huang, Pin-Yu Chen, et al. 2024. Justice or prejudice? quantifying biases in llm-as-a-judge. arXiv preprint arXiv:2410.02736.
@@
  • (43) Chujie Zheng, Hao Zhou, Fandong Meng, Jie Zhou, and Minlie Huang. Large language models are not robust multiple choice selectors, 2024. URL https://arxiv. org/abs/2309.03882.
  • Zheng et al. (2023) Mengxin Zheng, Qian Lou, and Lei Jiang. 2023. Trojvit: Trojan insertion in vision transformers. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pages 4025–4034.